OpenStudy (pnathaniel56):

what is 77/3 divided by 3/8

OpenStudy (1davey29):

In order to divide by a fraction, flip the second fraction and multiply. Ex) a/b divided by c/d becomes a/b times c/d. In your case, you have (77/3)/(3/8) which becomes (77/3)(8/3)=((7*8)/(3*3))=616/9=68+(5/9)

OpenStudy (triciaal):